OpenStudy (anonymous):

Solve: x − 5 = 10 A.15 B.10 C. 5 D. 0

OpenStudy (barrycarter):

Hint: Add 5 to both sides

OpenStudy (anonymous):

B

OpenStudy (zzr0ck3r):

x-5+5 = 10+5

OpenStudy (anonymous):

i am sure

OpenStudy (zzr0ck3r):

@katragaddasaichandra dont just give answers please.

OpenStudy (anonymous):

B

OpenStudy (anonymous):

I'ts Letter A.

OpenStudy (anonymous):

medal?

OpenStudy (anonymous):

you have to add 5 to both sides to get x by itself

OpenStudy (anonymous):

@zzr0ck3r hey

OpenStudy (zzr0ck3r):

I know you wanted to get the metal so you answered first, but that is no reason to call me stupid.

OpenStudy (anonymous):

do u look my explaination @zzr0ck3r

OpenStudy (zzr0ck3r):

just saying the rules m8, sorry. hope you get that medal.

OpenStudy (anonymous):

sorry @zzr0ck3r i am very sorry i am apologize

OpenStudy (anonymous):

Cuz 15 - 5 = 10 like 10 + 5 = 15

OpenStudy (anonymous):

and i will give you a hint as well, the answer is not b

OpenStudy (anonymous):

its A?

OpenStudy (anonymous):

guys no fighting